Operations Report – March 11th

Cerise, Crew 206 HSO.

Crew 206 Operations Report 11 – MAR – 2019

SOL: 15

Name of person filing report: Gaspard Thieulin

Non-nominal systems: Generator

Notes on non-nominal systems: Radiator hoses and
alternator need replacement on generator. A contractor has
been arranged to make these repairs within a few days.

Generator (hours run): 12,5 hours

Solar— SOC % – Turned off at 100%, Turned on at 78% (2000)

Diesel Reading – 60% Propane
Reading – 76%

Ethanol Free Gasoline – 2 gallons

Water (Auxillary tank) – Not in use

Water (Static tank) – 350 gallons

Water differential (static tank) – 13 gallons

Gallons transferred: Not Applicable

Water in GreenHab – 216 gallons

Static to Loft Pump used – Yes

Water Meter: 0141734,1 gallons

Toilet tank emptied: Yes

Deimos rover used: No, still not functional

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: N/A

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours: Not Applicable

Beginning charge: Not Applicable

Ending charge: Not Applicable

Currently charging: Not Applicable

Opportunity rover used: Breaking system being repared.

Hours: 53,1 hours

Beginning charge: 100%

Ending charge: 100%

Currently charging: No

Curiosity rover used: Used

Hours: 78.2 hours

Beginning charge: 100%

Ending charge: 97%

Currently charging: Yes

Spirit rover used: Used

Hours: 81,3 hours

Beginning charge: 100%

Ending charge: 93%

Currently charging: Yes

Notes on rovers: Opportunity’s brake fluid is empty, it looks like the
pipeline has a leak. Atila is working on it, more informations will follow.

ATV’s Used: N/A

Reason for use: Not used

Oil Added? N/A

ATV Fuel Used: N/A

# Hours the ATVs were Used today: None

Notes on ATVs: None

Hab Car used and why, where? Used by Staff.

Staff Crew Car used and why, where? Off Campus

General notes and comments: Nothing to report.

Summary of internet: Nothing to report.

Summary of suits and radios: Nothing to report.

Summary of Hab operations: Nothing to report.

Summary of Green Hab operations: Nothing to report.

Summary of Science Dome operations: Nothing to report.

Summary of RAM operations: Nothing to report.

Summary of any observatory issues: Nothing to report.

Summary of health and safety issues: Nothing to report

Questions, concerns and requests to Mission Support: Nothing more to report.

Journalist Report – March 11th

Cerise, Crew 206 HSO.

Crew 206 – 03/11/2019

Sol 15

Author : Benjamin Auzou, Journalist

"Trial and Error"

We installed the LOAC back to its location, with the weather station,
and a strongest attach method. The EVA was a remake of the second EVA of
the mission : same team, same goal.
After an afternoon in the Hab, the VR experiment took me to wonderful
landscapes and mountains of China. That was a visual refresh for me !
Today, Aurélien also received his first astronomy results since a few
days thanks to the end of the wind episode : a nice picture of the Pleiads.

This mission confirms that in space, long preparation is needed, but
that it’s impossible to predict all the events. The road to space is
full of difficulties, of obstacles, of errors. And problems are always
easier to solve when they’re detected before leaving the Earth.
We faced issues with the Teleop experiment, which wasn’t lucky at all
this year : a lost (or stolen) eye tracker at the airport, connection
issues on the rover, a broken computer charger. The weather was hard
with our outside experiments (LOAC and Weather station).
We prepared ourselves during one year for this mission but reality is
very different from what we thought. We had to adapt everything in order
to continue the mission and get results for science. We repaired the
weather station to put it back in place, we make changes on protocols to
save the physic version of Teleop. And in the end, no experiment has
been canceled !

The first step on Mars will be the result of years of work of peoples :
researchers, engineers, scientists who dedicated their life to improve
science and technology and help humanity to reach another planet. These
persons are the one that faced fails and that tried the most to make
discoveries, while the rest of the world will only see the final result
and success. Errors are not the end, but more steps to find the truth.

EVA Report – March 11th

Crew 206 EVA Report 11-MAR-2019
EVA #11

Author: Benjamin Auzou

Purpose of EVA:
After the wind of past days, the LOAC is back in the Hab for inside
measurements but we need to redeploy it outside for the last week. We
updated the water protection for these rainy days.

Start time: 09:59 am

End time: 11:04 am

Narrative:
This EVA was , a replay of our EVA#2 of 25th February. The same team and
the same goal.
After our engineering checks (water and rovers), we took Spirit and
Curiosity to drive to the LOAC location. The sky was blue and the sun
was shining on the martian landscape, so the EVA was quiet on the side
of weather.
We managed to install the station and attach it to the ground with pegs
and rocks. It took less time than on EVA#2 and in less than 45 minutes
the weather station and the LOAC were ready to restart their science.
We arrived at the station and entered the airlock at 1059. A successful
one our long EVA !

Destination: LOAC and Weather Station location

Coordinates: 518500/425700 (LOAC and Weather Station)

EVA Participants: Benjamin Auzou (COMMS, EVA leader), Gaspard Thieulin
(ENG), Cerise Cuny (HSO), Jérémy Auclair (CMD).

Roads and routs per MDRS Map: Entrance Road, Cow Dung Road

Mode of travel: Walking and Driving

Report Astronomy- March 11th

Astronomy Report, Sol 15

Name: Aurélien Mure
Crew: 206
Date: 03/11/2019

MUSK OBSERVARORY

Features observed : None

Problem encountered : I did two sessions of solar observation this
morning, no feature visible, even with the solar storm hitting us today.

ROBOTIC OBSERVATORY

Object observed : M45, the pleiades

Problems encountered : MaxImDL has severe bugs so I try to do image
processing with the other softwares, I plan various observations for
tonight to finish the plan I had for my mission. I attach the
astrophotography I did of M45.

GreenHab Report – March 11th

Crew 206 Green Hab Report – 11-MAR-2019

GreenHab Officer: Norbert Pouzin

Environmental control: Heating and Cooling w/ ambient air (8 hrs 20 min)

Shade Cloth (40%)

Average temperatures:

Low: 17.8°C

High: 30.3°C

Hours of supplemental light: 0 Hours

Daily water usage for crops: 5.6 gallons

Water in Blue Tank: 216.0 gallons

Crops watered at 0910 hours (2.2 gallons) and 1830 hours (3.2 gallons):

TIME NOTED TEMPERATURE(° C ) / HUMIDITY(%)

1) 09.10 AM 22.4° C / 29 %
2) 11.30 AM 24.6° C / 20 %
3) 15.20 PM 25.5° C / 20 %
4) 18.30 PM 19.5° C / 23 %

Changes to crops: Some lettuces and savory mix, planted on March 08, are
starting to sprout.

Narrative: This morning I harvested the last Savory mix that were in the
gardener, for lunch. In the afternoon, I continued my experiment, and
there are real signs of the influence of music, as a resistance
mechanism to radiations !
I opened the door from 9h10 to 17h30.
I harvested peas (12 peas), and soaked some pea seeds.

Harvest: 53 g of Savory mix, 2 g of peas

Support/supplies needed: None.

Sol Summary – March 11th

Crew 206 March 11th 2019
Sol 15

Summary Title : Weather Station round 2

Author : Jeremy Auclair

Mission Status : All nominal

Sol Activity Summary : We woke up still tired this morning, but
motivated to finish the mission as best we could. We continued filming a
lot of different scenes all day, this kind of changes our everyday
experience. This morning’s EVA was a success : we installed the weather
station and the LOAC again, but we anchored the base much more strongly
to the ground, it seems perfectly aligned, and we are getting the data.
This has also ended the 4 days of interior measurements, we are eager to
see the results of air quality we are breathing. This afternoon we
continued the TELEOP sessions, everything is running fine and we are not
encountering issues anymore. It feels good to finally have a working
set-sup after so much work. It’s also our last VR week, we are seeing
some pretty footage for the last time.

Look ahead plan : Tomorrow we will change the LOAC battery to make sure
it doesn’t turn off, we will try to visit white rock canyon and we will
pursue our experiments.

Anomalies in work : None

Weather : Quite sunny until 4:30 pm, Cloudy after that

Crew physical status : Fine

EVA : LOAC and weather station installation

Reports to be filed : GreenHab, Journalist, Operation, EVA Request, EVA
Report

Support Requested : None